40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Jul 22 20:51:13 2011, frank, tara, DailyProgress, optic, cleaning opto mechanical parts IMG_1862.JPGIMG_1863.JPG
    Reply  Tue Jul 26 18:42:16 2011, raphael, tara, DailyProgress, optic, cleaning opto mechanical parts IMG_1866.JPGIMG_1867.JPG
       Reply  Thu Jul 28 01:54:21 2011, frank, tara, DailyProgress, optic, cleaning opto mechanical parts 2011_07_28.pngIMG_1869.JPGIMG_1874.JPGIMG_1875.JPG
          Reply  Fri Jul 29 00:42:19 2011, frank, tara, DailyProgress, optic, cleaning opto mechanical parts IMG_1880.JPGIMG_1883.JPGIMG_1889.JPGIMG_1886.JPGIMG_1878.JPG
             Reply  Tue Aug 2 23:58:40 2011, tara, DailyProgress, optic, New RCAV setup is locked IMG_1911.JPGrcav_2011_08_02.jpg
                Reply  Tue Aug 9 00:39:30 2011, tara, DailyProgress, optic, RCAV modematching optimized 
                   Reply  Mon Aug 15 23:13:53 2011, tara, DailyProgress, optic, RCAV modematching optimized 2011_08_15.png
                      Reply  Wed Aug 17 20:41:13 2011, tara, DailyProgress, optic, ACAV path is up layout_2011_08_17.jpg
Message ID: 640     Entry time: Tue Jul 26 18:42:16 2011     In reply to: 636     Reply to this: 644
Author: raphael, tara 
Type: DailyProgress 
Category: optic 
Subject: cleaning opto mechanical parts 

Today we removed the optics behind the PMC, ACAV, and cleaned the table.

  •   Optic mounts and posts are cleaned by pressurized air, and kept in a plastic box.
  •   Lens and mirrors are kept in optic cases.
  •   ACAV is moved to the end of the table, ion pump is unplugged.
  • Table is cleaned with methanol, but some grease( under acav) is still on it.

NOTE: I just realized that the HEPA filter above the table close to the entrance ( the one that has the laser) is unplugged.

I could not find any available outlet to plug it back yet. We should turn it on soon.

IMG_1867.JPG

 

 

IMG_1866.JPG

ELOG V3.1.3-